OpenStudy (anonymous):

Mike wants to make meatloaf. His recipe uses a total of 6 pounds of meat. If he uses a 3 to 1 ratio of beef to pork, how much pork will he use? Answer needs to be as a mixed number in simplest terms.

OpenStudy (anonymous):

he uses 2 lbs of pork

OpenStudy (anonymous):

he would use 2 pounds of pork

OpenStudy (anonymous):

It was incorrect, why was it ;o.

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

It was incorrect it said that 2 was incorrect ;o

OpenStudy (anonymous):

oh okay let me solve it again

OpenStudy (anonymous):

We used 8 pounds.

OpenStudy (anonymous):

Mike will have 3 parts of beef and 1 part of pork. That totals 4 parts. He will need to divide 6 pounds ÷ 4 parts = 1.5 pounds . Now Mike has to multiply 1.5 pounds * 3 to get the amount of beef. 3 parts of beef * 1.5 pounds per part = 4.5 pounds of beef. 1 part of port * 1.5 pounds per part = 1.5 pounds of pork. 4.5 + 1.5 = 6 pounds 4.5 = 4 1/2 1.5 = 1 /2

OpenStudy (anonymous):

that seems right :)